OpenStudy (anonymous):

simplify (5 square root 40)

OpenStudy (anonymous):

5 sqrt 40 5 sqrt 4 * 10 5 * 2 sqrt 10 10 sqrt 10

OpenStudy (agreene):

\[10\sqrt{10}\]

OpenStudy (anonymous):

10sqrt(10)

OpenStudy (anonymous):

\[10\sqrt{10}\]

OpenStudy (anonymous):

sqrt(a)*sqrt(b) = sqrt(a*b)

OpenStudy (anonymous):

man you guys are great, THANKS!

OpenStudy (anonymous):

ok what about this

OpenStudy (callisto):

\[5\sqrt[3]{96} = 5\sqrt[3]{2\times2\times2\times2\times2\times3} = 5\times 2\sqrt[3]{2\times2\times3} = 5\times 2\sqrt[3]{12} \]

OpenStudy (muhammad_nauman_umair):

10sqrt(12)

OpenStudy (anonymous):

\[10\sqrt{12}\]

OpenStudy (anonymous):

do i foil this???